OpenStudy (anonymous):

I GIVE MEDALS AND I RLYY NEED HELP?!? PLZ JUST TRY IT EVEN IF U CANT FINISH IT WILL STILL HELP ME OUT :) THANKS sin6x= 2sinxcosx(3-4sin^2 x)(1-4sin^2 x)

OpenStudy (michele_laino):

please use the Euler formula, namely: \[{e^{ix}} = \cos x + i\sin x\] so we can write: \[{e^i}^{6x} = \left( {\cos \left( {6x} \right) + i\sin \left( {6x} \right)} \right) = {\left( {\cos x + i\sin x} \right)^6}\]

OpenStudy (michele_laino):

please note that you have to develop the 6-th power

OpenStudy (anonymous):

we havent learned that formula - i think its supposed to use double angle formulas

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

then you have to evaluate \[\cos \left( {3x} \right) = \cos \left( {2x + x} \right) = ...?\] first, and then you have to evaluate \[\cos \left( {6x} \right) = \cos \left( {2 \cdot 3x} \right) = 2{\left( {\cos \left( {3x} \right)} \right)^2} - 1\]

OpenStudy (anonymous):

im confused on why you are using cos (6x) because its for sin (6x)

OpenStudy (anonymous):

my teacher started it by breaking the sin(6x) into sin(3x+3x) but they never finished it. could u try walking through the steps that way please?

OpenStudy (michele_laino):

oops.. you are right!

OpenStudy (michele_laino):

Please I re-write your identities

OpenStudy (anonymous):

what do u mean rewrite them?

OpenStudy (michele_laino):

\[\begin{gathered} \cos \left( {3x} \right) = \cos \left( {2x + x} \right) = ...? \hfill \\ \sin \left( {3x} \right) = \sin (2x + x) = ...? \hfill \\ \sin \left( {6x} \right) = \sin \left( {2 \cdot 3x} \right) = 2\sin \left( {3x} \right)\cos \left( {3x} \right) = ...? \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

please apply the above identities, in order to find your anser

OpenStudy (anonymous):

okay so then would u break up 2sin(3x)cos(3x) into 2sin(2x+x)cos(2x+x) right?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

2(sin2xcosx+cos2xsinx)(cos2xcosx-sin2xsinx) ?

OpenStudy (anonymous):

what would u do after that?

OpenStudy (michele_laino):

please you have to apply the subsequent identities: \[\begin{gathered} \sin \left( {2x} \right) = 2\sin x\cos x \hfill \\ \cos \left( {2x} \right) = {\left( {\cos x} \right)^2} - {\left( {\sin x} \right)^2} \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

i know but i don't understand how it's going to get to the (3-4sin^2x)(1-4sin^2x) part..

OpenStudy (anonymous):

@directrix could u help me out?

OpenStudy (anonymous):

@michele_laino im rly confused could u just write out what the next line would look like... i feel like maybe its supposed to use the cos2x=1-sin^2x and 2cos^2x-1 instead of cos^2x-sin^2x because how else would u get the second half of the identity...?

OpenStudy (michele_laino):

please wait, I try

OpenStudy (anonymous):

okay thank you

OpenStudy (michele_laino):

\[\begin{gathered} 2\left[ {2\sin x\cos x\cos x + \left( {1 - 2{{\left( {\sin x} \right)}^2}} \right)\sin x} \right] \cdot \left[ {\left( {1 - 2{{\left( {\sin x} \right)}^2}} \right)\cos x - 2\sin x\cos x\cos x} \right] = \hfill \\ = 2\left[ {2\sin x\cos x} \right] \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

thank you so much! but is there a missing cos in the first term and a missing sin in the last term? if so would u just multiply those in

OpenStudy (anonymous):

nevermind!!

OpenStudy (michele_laino):

\[\begin{gathered} = 2\left[ {2\sin x{{\left( {\cos x} \right)}^2} + \sin x - 2{{\left( {\sin x} \right)}^3}} \right] \cdot \hfill \\ \cdot \left[ {\cos x - 2{{\left( {\sin x} \right)}^2}\cos x - 2{{\left( {\sin x} \right)}^2}\cos x} \right] = \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

\[\begin{gathered} = 2\left[ {2\sin x\left( {1 - {{\left( {\sin x} \right)}^2}} \right) + \sin x - 2{{\left( {\sin x} \right)}^3}} \right] \cdot \hfill \\ \cdot \left[ {\cos x - 4{{\left( {\sin x} \right)}^2}\cos x} \right] \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

and now it is easy to get your answer, please complete the computation

OpenStudy (anonymous):

thank you a lot you really helped! ;)

OpenStudy (michele_laino):

thank you! :)